Question 3 A ball is thrown upward at a speed vo at an angle of 48.0 above the horizontal. It reaches a maximum height of 8.5 m. How high would this ball go if it were thrown straight upward at speed vo?
Answer #1

here, for the speed v0

Height = (v0 * sin(48 degree))^2/(2 * g)

8.5 = (v0 * sin(48 degree))^2/(2 * 9.8)

v0 = 17.4 m/s

if the ball is thrown upwards

Height reached by ball = v0^2/(2 * 9.8)

Height reached by ball = 17.4^2/(2 * 9.8)

Height reached by ball = 15.4 m

the height reached by the ball is 15.4 m